OpenStudy (anonymous):

Given the points (2, 3) and (6, 8), write the point-slope form of the equation. y - 3 = x - 2 y - 3 = 5/4 (x - 2) y - 8 = 4/5 (x - 6)

OpenStudy (anonymous):

i think its the second on ebut am not sure

OpenStudy (mathlegend):

First, find the slope.

OpenStudy (mathlegend):

y2-y1/x2-x1

OpenStudy (mathlegend):

Can you substitute in the values?

OpenStudy (anonymous):

5/4

OpenStudy (mathlegend):

Good

OpenStudy (mathlegend):

Now use the point slope form formula. y-y1=m(x-x1)

OpenStudy (anonymous):

whats the next step?

OpenStudy (mathlegend):

Plug in those values

OpenStudy (anonymous):

um y=mx+b so y=5/4x+b

OpenStudy (mathlegend):

No.... just plug those values into the point slope form and what did you get?

OpenStudy (anonymous):

point slope form?

OpenStudy (mathlegend):

y-y1=m(x-x1)

OpenStudy (anonymous):

m=y-y1/x=x1?

OpenStudy (mathlegend):

y-y1=m(x-x1) y-3=5/4(x-2) All I did was substitute in things. Do you understand @EmmaH

OpenStudy (anonymous):

you only used the first pair? o

OpenStudy (mathlegend):

Yes, because it is... y-Y1=M(x-X1)

OpenStudy (mathlegend):

It does not ask for x2 and y2

OpenStudy (anonymous):

oh ok so y-3=5/4(x-2) is the answer?

OpenStudy (mathlegend):

Yes, that is the point slope form.

OpenStudy (anonymous):

ok thanks!=D